How do I write a linear function f with the values f(−2)=3 and f(5)=7?
Answer by Alan3354(69443)

How do I write a linear function f with the values f(-2)=3 and f(5)=7?
----
Use the 2 points, (-2,3) and (5,7)
Slope m = diffy/diffx = 4/7
-----
y-y1 = m*(x-x1)
y-7 = (4/7)*(x-5)
y = 4x/7 + 29/7
f(x) = 4x/7 + 29/7